# Geometry-Controlled Work Extraction in a Non-Markovian Quantum Battery

### v1

@@ Abstract
Figure 4: Geometry and espectral width dependence of quantum battery performance for $J=-0.005\,\omega_{0}$. (a) Maximum internal energy $\Delta E_{B}^{\max}$, (b) maximum charging power $\mathcal{P}_{B}^{\max}$, and (c) maximum ergotropy $\mathcal{W}_{B}^{\max}$ as functions of the geometric parameter $d/\lambda_{0}$ and normalized reservoir spectral width $\lambda/g$. Optimal performance occurs at small $\lambda/g$ and geometries supporting constructive collective interference.
Fig. 4 summarizes the geometry spectral width dependence of the quantum battery performance. Panel (a) shows the maximum internal energy $\Delta E_{B}^{\max}$, panel (b) the maximum charging power $\mathcal{P}_{B}^{\max}$, and panel (c) the maximum ergotropy $\mathcal{W}_{B}^{\max}$ as functions of the geometric parameter $d/\lambda_{0}$ and the normalized reservoir spectral width $\lambda/g$. In all three panels, the battery performance is strongly enhanced in the small-spectral width regime ($\lambda/g\ll 1$), where reservoir memory effects are significant. As $\lambda/g$ increases, both the maximum stored energy and the extractable work decrease monotonically, indicating the progressive suppression of coherent energy exchange in the Markovian limit. This behavior reflects the reduction of information backflow and the transition toward irreversible dissipation. Along the geometric axis, a clear periodic modulation appears in all quantities. The maxima occur at specific separations corresponding to constructive collective interference, where both symmetric and antisymmetric channels contribute effectively to the dynamics. In contrast, reduced performance is observed near configurations associated with partial destructive interference. Importantly, the locations of optimal geometry are consistent across internal energy, charging power, and ergotropy, demonstrating that geometry acts as a universal control parameter for both energy storage and work extraction. While $\Delta E_{B}^{\max}$ and $\mathcal{W}_{B}^{\max}$ exhibit similar spatial structures, the charging power map shows a sharper sensitivity to spectral width, indicating that fast charging is more strongly affected by reservoir-induced decoherence than the total stored energy. Overall, Fig. 4 demonstrates that optimal quantum battery performance is achieved in the combined regime of strong collective interference (appropriate geometry) and long reservoir memory (small $\lambda/g$), highlighting the cooperative role of spatial phase engineering and non-Markovian dynamics.
I Conclusion
@@
Appendix A Exact Single-Excitation Dynamics in a Lorentzian Reservoir
In this model, we assume that the quantum battery is initially empty, while the charger is prepared in its excited state. This corresponds to choosing the initial state with the conditions $\alpha_{1}(0)=1$$\alpha_{2}(0)=0$ and $\beta_{k}(0)=0$ in Eq. (11). By substituting the total state of the system at time $t$, together with the total Hamiltonian, into the Schrödinger equation, one obtains the following set of coupled differential equations
$\dot{\alpha}_{1}(t)$ $=-i\omega_{0}\alpha_{1}(t)-iJ\alpha_{2}(t)-i\sum_{k}g_{k}^{(1)}\,\beta_{k}(t),$ (19)
$\dot{\alpha}_{2}(t)$ $=-i\omega_{0}\alpha_{2}(t)-iJ\alpha_{1}(t)-i\sum_{k}g_{k}^{(2)}\,\beta_{k}(t),$ (20)
$\dot{\beta}_{k}(t)$ $=-i\omega_{k}\beta_{k}(t)-i\Big[g_{k}^{(1)*}\alpha_{1}(t)+g_{k}^{(2)*}\alpha_{2}(t)\Big].$ (21)
Solving Eq. (21) with $\beta_{k}(0)=0$ yields the formal bath amplitude
$\beta_{k}(t)=-i\int_{0}^{t}d\tau\,e^{-i\omega_{k}(t-\tau)}\Big[g_{k}^{(1)*}\alpha_{1}(\tau)+g_{k}^{(2)*}\alpha_{2}(\tau)\Big].$ (22)
Substituting (22) into (19)–(20) eliminates the bath and produces memory integrals. Using the standard phase approximation for spatially separated emitters,
$g_{k}^{(1)}\simeq g_{k}e^{+ik_{0}d},\qquad g_{k}^{(2)}\simeq g_{k}e^{-ik_{0}d},$ (23)
one obtains the integro-differential system
$\small\dot{\alpha}_{1}(t)$ $=$ $-i\omega_{0}\alpha_{1}(t)-iJ\alpha_{2}(t)$ (24)
- $\int_{0}^{t}d\tau\,K(t-\tau)\Big[\alpha_{1}(\tau)+e^{+i2k_{0}d}\alpha_{2}(\tau)\Big],$
$\dot{\alpha}_{2}(t)$ $=$ $-i\omega_{0}\alpha_{2}(t)-iJ\alpha_{1}(t)$ (25)
- $\int_{0}^{t}d\tau\,K(t-\tau)\Big[\alpha_{2}(\tau)+e^{-i2k_{0}d}\alpha_{1}(\tau)\Big],$
where the bath correlation kernel is
$K(t)\equiv\int_{0}^{\infty}d\omega\,J(\omega)\,e^{-i\omega t},\quad(t\geq 0).$ (26)
By considering the collective amplitudes defined in the main text and using $e^{\pm i2k_{0}d}=\cos(2k_{0}d)\pm i\sin(2k_{0}d)$, the above integro-differential equations can be rewritten in the following form
$\dot{s}(t)$ $=-i(\omega_{0}+J)\,s(t)$
$\quad-\int_{0}^{t}d\tau\,K(t-\tau)\zeta(\tau),$ (27)
$\dot{a}(t)$ $=-i(\omega_{0}-J)\,a(t)$
$\quad-\int_{0}^{t}d\tau\,K(t-\tau)\xi(\tau),$ (28)
where $\zeta(\tau)=\Big[2\cos^{2}(k_{0}d)\,s(\tau)+i\sin(2k_{0}d)\,a(\tau)\Big]$ and $\xi(\tau)=\Big[2\sin^{2}(k_{0}d)\,a(\tau)-i\sin(2k_{0}d)\,s(\tau)\Big]$. We now assume a Lorentzian spectral density centered at $\omega_{0}$,
$J_{L}(\omega)=\frac{\gamma\lambda^{2}}{(\omega-\omega_{0})^{2}+\lambda^{2}},$ (29)
with coupling scale $\gamma$ and linewidth $\lambda$ (memory time $\tau_{M}\sim\lambda^{-1}$). Evaluating (26) for (29) yields an exponential kernel
$K(t)=g^{2}e^{-\tilde{\lambda}t}\,\Theta(t),\quad g=\sqrt{\frac{\gamma\lambda}{2}},\quad\tilde{\lambda}=\lambda-i\omega_{0},$ (30)
where $\Theta(t)$ is the Heaviside function. Because $K(t)$ is exponential, the convolutions in Eqs. (27)–(28) can be made time-local by introducing two auxiliary amplitudes
$\beta(t)$ $=$ $ig\int_{0}^{t}d\tau\,e^{-\tilde{\lambda}(t-\tau)}\,s(\tau),$
$\alpha(t)$ $=$ $ig\int_{0}^{t}d\tau\,e^{-\tilde{\lambda}(t-\tau)}\,a(\tau).$ (31)
So, the dynamics become an exact closed set of first-order ODEs for the four amplitudes $(s(t),a(t),\beta(t),\alpha(t))$:
$\dot{s}(t)$ $=$ $-i(\omega_{0}+J)\,s(t)$ (32)
+ $+\,2ig\cos^{2}(k_{0}d)\,\beta(t)-g\sin(2k_{0}d)\,\alpha(t),$
$\dot{a}(t)$ $=$ $-i(\omega_{0}-J)\,a(t)$ (33)
+ $+\,2ig\sin^{2}(k_{0}d)\,\alpha(t)+g\sin(2k_{0}d)\,\beta(t),$
$\dot{\beta}(t)$ $=$ $ig\,s(t)-\tilde{\lambda}\,\beta(t),$ (34)
$\dot{\alpha}(t)$ $=$ $ig\,a(t)-\tilde{\lambda}\,\alpha(t).$ (35)
with initial conditions
$s(0)=\frac{1}{\sqrt{2}},\qquad a(0)=\frac{1}{\sqrt{2}},\qquad\beta(0)=0,\qquad\alpha(0)=0.$ (36)
The exact equations (32)–(35) can be written in compact vector form. Define the state vector
$\mathbf{X}(t)=\begin{pmatrix}s(t)\\ a(t)\\ \beta(t)\\ \alpha(t)\end{pmatrix},\qquad\mathbf{X}(0)=\begin{pmatrix}\frac{1}{\sqrt{2}}\\ \frac{1}{\sqrt{2}}\\ 0\\ 0\end{pmatrix}$. (37)
Then the dynamics are governed by the linear first-order system
$\dot{\mathbf{X}}(t)=M(d)\,\mathbf{X}(t),$ (38)
where the distance-dependent evolution matrix is
$\small M(d)=\begin{pmatrix}-i(\omega_{0}+J)&0&2ig\cos^{2}(k_{0}d)&-g\sin(2k_{0}d)\\ 0&-i(\omega_{0}-J)&g\sin(2k_{0}d)&2ig\sin^{2}(k_{0}d)\\ ig&0&-\tilde{\lambda}&0\\ 0&ig&0&-\tilde{\lambda}\end{pmatrix}$, (39)
Formal solution.
Because Eq. (38) is linear and time-independent, its formal solution is
$\mathbf{X}(t)=e^{M(d)t}\,\mathbf{X}(0),$ (40)
where the matrix exponential is defined by the convergent series
$e^{M(d)t}=\sum_{n=0}^{\infty}\frac{[M(d)t]^{n}}{n!}.$ (41)
Eigenmode decomposition.
If $M(d)$ is diagonalizable, it admits the similarity decomposition
$M(d)=P(d)\,\Lambda(d)\,P^{-1}(d),$ (42)
where
$\Lambda(d)=\mathrm{diag}(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ (43)
contains the complex eigenvalues $\mu_{i}(d)$ and the columns of $P(d)$ are the corresponding eigenvectors.
The formal solution then becomes
$\mathbf{X}(t)=P(d)\,e^{\Lambda(d)t}\,P^{-1}(d)\,\mathbf{X}(0),$ (44)
or equivalently
$\mathbf{X}(t)=\sum_{i=1}^{4}c_{i}\,\mathbf{v}_{i}\,e^{\mu_{i}t},$ (45)
where
$c_{i}=\big[P^{-1}(d)\mathbf{X}(0)\big]_{i}$ (46)
and $\mathbf{v}_{i}$ are the eigenvectors of $M(d)$.
So, using the eigenmode expansion (45), one obtains the explicit general solution
$\alpha_{1}(t)=\frac{1}{\sqrt{2}}\sum_{i=1}^{4}c_{i}\,(v_{i,s}+v_{i,a})\,e^{\mu_{i}t},$ (47)
$\alpha_{2}(t)=\frac{1}{\sqrt{2}}\sum_{i=1}^{4}c_{i}\,(v_{i,s}-v_{i,a})\,e^{\mu_{i}t},$ (48)
where $v_{i,s}$ and $v_{i,a}$ denote the symmetric and antisymmetric components of eigenvector $\mathbf{v}_{i}$.

### v2

@@ III Physical model
Figure 4: (a) Maximum internal energy $\Delta E_{B}^{\max}$, (b) maximum charging power $\mathcal{P}_{B}^{\max}$, and (c) maximum ergotropy $\mathcal{W}_{B}^{\max}$ as functions of the geometric parameter $l/\lambda_{0}$ and $\lambda/\gamma$. Optimal performance occurs at small $\lambda/\gamma$ and geometries supporting constructive collective interference.
Fig. 4 summarizes the geometry spectral width dependence of the quantum battery performance. Panel (a) shows the maximum internal energy $\Delta E_{B}^{\max}$, panel (b) the maximum charging power $\mathcal{P}_{B}^{\max}$, and panel (c) the maximum ergotropy $\mathcal{W}_{B}^{\max}$ as functions of the geometric parameter $l/\lambda_{0}$ and $\lambda/\gamma$. In all three panels, the battery performance is strongly enhanced in the small-spectral width regime ($\lambda/\gamma\ll 1$), where reservoir memory effects are significant. As $\lambda/\gamma$ increases, both the maximum stored energy and the extractable work decrease monotonically, indicating the progressive suppression of coherent energy exchange in the Markovian limit. This behavior reflects the reduction of information backflow and the transition toward irreversible dissipation. Along the geometric axis, a clear periodic modulation appears in all quantities. The maxima occur at specific separations corresponding to constructive collective interference, where both symmetric and antisymmetric channels contribute effectively to the dynamics. In contrast, reduced performance is observed near configurations associated with partial destructive interference. Importantly, the locations of optimal geometry are consistent across internal energy, charging power, and ergotropy, demonstrating that geometry acts as a universal control parameter for both energy storage and work extraction. While $\Delta E_{B}^{\max}$ and $\mathcal{W}_{B}^{\max}$ exhibit similar spatial structures, the charging power map shows a sharper sensitivity to spectral width, indicating that fast charging is more strongly affected by reservoir-induced decoherence than the total stored energy. Overall, Fig. 4 demonstrates that optimal quantum battery performance is achieved in the combined regime of strong collective interference (appropriate geometry) and long reservoir memory (small $\lambda/\gamma$), highlighting the cooperative role of spatial phase engineering and non-Markovian dynamics.
IV Conclusion